OpenStudy (anonymous):

Can someone please help me with this. I don't understand how to do this. There are 2 more problem similar to this but if someone helps me I can finish those as well. Use the formula: F +Vn =Pn; where F is fixed cost, V is variable cost, P is selling price, and "n" is the number of units produced. Assume a new product has a fixed cost of $900, a variable cost of $4.50 each, and a selling price of $5.50 each. (a) What number of units must be produced to "break even?

OpenStudy (texaschic101):

F + Vn = Pn F = 900 V = 4.50 P = 5.50 now we sub 900 + 4.5n = 5.5n -- subtract 5.5n from both sides 900 = 5.5n - 4.5n 900 = n lets check... 900 + 4.5(900) = 5.5(900) 900 + 4050 = 4950 4950 = 4950 (correct) To break even, 900 units must be produced

OpenStudy (texaschic101):

oops..typo...I meant to put " subtract 4.5n from both sides "

OpenStudy (anonymous):

so 900-4.5=5.5n-4.5n

OpenStudy (texaschic101):

no... 900 + 4.5n - 4.5n = 5.5n - 4.5n -- combine like terms 900 = 5.5n - 4.5n 900 = n

OpenStudy (texaschic101):

do you understand ?

OpenStudy (anonymous):

I'm starting to see what you mean.

OpenStudy (anonymous):

So for for this problem I would replace 900 with 500 right? How many units must be sold to make a profit of $500?

OpenStudy (texaschic101):

Once we subbed in our info, we got this.. 900 + 4.5n = 5.5n -- we want to get the n's on one side, so I subtracted 4.5n from both sides, because what you do on one side of the equal sign, you have to do on the other side. I subtracted 4.5n from both sides 900 + 4.5n - 4.5n = 5.5n - 4.5n -- combine like terms 900 = n you understand now ?

OpenStudy (anonymous):

okay I see what you mean you had to make the 4.5 a negative to bring it over on both sides to cancel out.

OpenStudy (texaschic101):

900 units to break even.... 900 + 500 = 1400 900 + 4.5(1400) = 5.5(1400) 900 + 6300 = 7700 7200 = 7700 --- here, you would make a profit of $500 if you sold 1400 units

OpenStudy (anonymous):

so the n you substituted it for 1400 on both sides.

OpenStudy (anonymous):

How many units must be sold to average $0.25 profit per unit? $0.50 profit per unit? (hint: Profit per unit = P - V - F/n) so I did: 5.50-4.50-900/900=0

OpenStudy (anonymous):

Thanks for your help. I really appreciate you helping me understand this.

OpenStudy (texaschic101):

I am not sure about the profit per unit.....sorry, I can't really help with that

OpenStudy (anonymous):

That's okay. Thank you though.

OpenStudy (texaschic101):

your welcome :)